### v1

@@ 2.4 RQ4: Thresholds
As apparent from the figure, the implementation of a threshold has an immediate negative impact on quest completion. The purple lines follow the same trend as the grey (control) lines (dampened by not being easy tasks) until the threshold is implemented, and then plummet. One might hypothesize that bots and farmers would take a few epochs to update their scripts or acquire threshold tokens, but the impact appears sticky with no future uptick. Removing the threshold appears to return the quest to the control group’s trend line.
@@
Similarly, another quest, which had a threshold of approximately \$5 USD enforced at epoch 23, experienced a 75.8% drop in completions. After the threshold was removed at epoch 33, completions increased by 33.4%. These examples highlight the substantial impact of thresholds on user participation and the positive effects of removing them.
2.5 RQ5: Rewards with Monetary Value
@@
Figure 6: Five quests (by colour) that were offered over multiple epochs with a reward in non-monetary points (bars with no star) and also offered for one epoch with a token reward (bar with star). Grey line shows an average of task completion for the other quests over the same epochs. The charts shows that offering tokens (with a monetary value) as rewards boosts quest completion.
@@
Consider Figure 6 which depicts 5 quests by colour, each offered over epochs 43–49 (expect the yellow quest, which was first offered in epoch 46). For most epochs, the completion award was non-monetary (loyalty or experience) points, as we have analyzed in the previous section. However for one epoch (shown with a star), alongside the usual point rewards, specific quests were selected for a token reward (in the native token of the blockchain project) which had a small monetary value of  \$0.10 USD. The grey line shows an average completion rate of the other tasks offered across the same epochs.
@@
This was not a perfect experiment to determine the impact of monetary rewards because the rewards were capped at \$5000 USD, allowing a maximum of 50,000 users to claim rewards before they ran out. Further if two quests offered monetary rewards in the same epoch, they would compete against each other until the tokens were depleted (see epoch 46 in Figure 6).
@@
Despite these defects in the data, the results are insightful: even with a minimal reward, completion rates surged. The highest increase was over 12x for a quest typically completed by around 3,500 users. With the token reward, this quest (purple in Figure 6) saw 44,196 completions. Directly after the token reward for this quest was removed, the completion rate dropped back to its original levels before the reward was introduced.
@@
What this data reveals is that adding a token reward results in a spike in completions, but once the reward is removed, completion rates drop again. These spikes suggest that token rewards do not foster loyalty or habitual engagement but rather create a purely transactional relationship. Users appear to compare the cost to the reward and only engage when there is a clear opportunity for profit.
@@
In epoch 46, two quests with token rewards showed different outcomes. While the yellow quest, which had a low cost, saw a significant increase in completions, the orange quest, which had a high cost, experienced no increase at all. If users were engaging with the platform out of loyalty or mastery, they would have completed the orange quest as well as completing the yellow quest. However this did not happen: users who claimed the reward for the yellow quest did not complete the orange quest, leaving the reward for the orange quest unclaimed.
2.6 RQ6: Bots and farming
We currently lack data revealing how many task completions come from unique individuals versus repeat participants using multiple addresses (farming), whether by manual sybils or automated bots. Both human users and bots seem drawn to quests with higher monetary rewards, simpler tasks, and lower costs, making it difficult to distinguish them based on these factors alone.
Possible distinguishers.
The Fogg Behavior Model [3] was proposed to understand how well commercial products integrate into users’ daily routines, and we see it as relevant to the design of quests. The model is expressed as the pseudo-equation $B=MAP$ where a behaviour (B) requires simultaneous motivation (M), ability (A), and a prompt (P) (formerly known as a trigger). If any one of these components is missing or inadequate, establishing a specific behaviour is challenging or impossible.
Under this model, do non-transactional users (‘genuine users’) and farmers, bots, and scripts (‘bots’) exhibit the same behaviour (B) when the inputs (MAP) are the same? The answer seems to be no. Both are motivated (M) when the benefits outweigh the costs, but genuine users are largely unaffected when the benefits are small (cf. picking up a penny off the road). By contrasts, bot farms will immediately engage because they can use scale to accumulate a tiny profit from each quest into a material amount (cf. picking up thousands of pennies).
Bots are more responsive to prompts (P) (e.g., a new quest is available) because they are scripted to listen for them, whereas genuine users need rely on marketing, reading push notifications, or remembering to check the website. At an extreme, bots monitoring the blockchain’s mempool for profitable transactions might discover profitable quest completions without knowing anything about the quests themselves. Such bots are called general-purpose Maximal Extractable Value (MEV) bots [2] and they simply re-stage every transaction they see before the transaction is confirmed, and if they determine it will be profitable for them to originate, they will front-run the original transaction. The ability (A) to react quickly to prompts, and at the extreme, front-run other transactions, is another distinguishing feature of bots versus genuine users.
Finally, we observe that to complete a quest, an address needs to hold an adequate amount of currency and/or number of tokens to have the ability (A) to pay gas fees and any other fees associate with the quest. Genuine users are likely to hold more than the minimum amount and to have held it for some period of time. By contrast, farming operations have to divy the tokens across many addresses, meaning there is likely a correlation between these addresses being newly created and holding a minimal amount of capital.
Putting this together, we suggest three heuristics that may help distinguish genuine users from bots. There are likely exceptions to all of them (e.g., genuine humans running scripts) but we hypothesize that they would do better than random guessing at determining if a given address is a genuine user or a bot.
  • Heuristic 1: If a quest goes from gradually becomes profitable over time, bots will be sensitive to the exact moment it becomes marginally profitable and act.
  • Heuristic 2: If a quest requires a threshold (minimal amount of tokens), bots will be sensitive to ensuring the address meets but does not exceed the threshold amount.
  • Heuristic 3: If a (profitable) quest is announced (especially unexpectedly), bots will be the quickest to react.
Finally we emphasize that if these heuristics were deployed in a way that harmed bots (e.g., banning rather than passive measurement), bots can adjust their behaviour to evade detection, resulting in a ‘cat and mouse’ game between farmers and quest creators.
Quest design ideas.
In the absence of concrete measurements and based on these insights, we outline several ideas for collecting more robust data, and designing better quests that can help identify and quantify farming activity.
@@ 2.6 RQ6: Bots and farming
Building on heuristic 1, if costs (i.e., gas fees, service fees, buying a token or NFT as part of the quest, etc.) are slightly higher than rewards (i.e., tokens of monetary value received for quest completion), genuine users might be happy they are at least receiving a subsidy for a quest they would complete anyways, while bots will avoid unprofitable tasks. When benefits are slightly higher than costs, bots will engage. For many quest designs, the profitability of the quest could be due to external factors such as fluctuations in gas rates or in the value of the token received. This could create the conditions for a ‘natural experiment’ where a quest is offered over an epoch, but its profitability fluctuates between negative and positive. Activity spikes when it becomes positive could be largely attributed to bots.
A second idea corresponds to heuristic 2 and the use of thresholds (RQ5). If a quest can only be completed by addresses holding at least $X$ tokens, bot farms may need to move $X$ tokens between their addresses. In general, clustering algorithms can use heuristics on blockchain data to suggest addresses belonging to the same user, and inducing these kinds of transfers could help make the pattern more distinguishable. Farming can also be frustrated by requiring each address to have held $X$ tokens for a certain period of time (days or weeks). Of course, if a farm with $n$ bots holds more than $n\cdot X$ tokens, farms will be unaffected by the use of thresholds.
A final idea builds on heuristic 3. In the quest system we studied, a smart contract was randomly replenished with a fixed daily reward pool of \$500. Pre-enrolled participants with sufficient loyalty points could claim a nominal reward (\$0.01 or \$0.025, depending on tier) until the pool was depleted.
@@
Looking at Table 1, we observe a marked decrease in the time taken to deplete all rewards as the days progressed. This would be expected if participants developed a habit of claiming rewards at a consistent daily time; however, column 2 indicates that refill times varied significantly, making it difficult for users to predict when rewards would be replenished. The key observation lies in the number of claims recorded within the first 2 minutes of refilling the contract, demonstrating the dominance of automated systems.
On Day 7, for instance, despite a one-hour delay in the refill time, approximately 1,500 claims were made within 2 minutes. The following day, a more drastic delay of 3 hours saw claims drop to 300, likely due to the implementation of rate-limiting and a mandatory wait period between claims. Notably, this led to a temporary spike in the time required to deplete all rewards, which rose to 2h41m. By Day 9, however, the refill occurred 4.5 hours earlier, and bots had clearly adapted, with 1,500 claims occurring within 2 minutes and the time to deplete rewards dropping to 27 minutes. Finally, following the introduction of an automated daily refill schedule on Days 10 and 11, the system reached its operational minimum, with rewards being depleted in under 30 minutes and claims stabilizing at  1,500 within the first 2 minutes.
@@
By studying claims made very quickly or even launching a quest such as ‘make a claim within the first 2 minutes after rewards were added,’ it is likely that observed completions could be attributed to bots. Even without conducting in-depth analysis, many teams (based on personal conservations and social media) sense that their platform has attracted a significant number of bots. In some instances, these teams recognize that their applications or quest systems have effectively become ‘farming machines’ for bots. However, as discussed in the next section, awareness of such bot activity does not necessarily compel teams to intervene; in certain scenarios, they may even be incentivized to sustain it.
3 Stakeholder Analysis
@@
A deeper question is what stakeholders actually prioritize. Is it the elimination of bots, metric boosts, revenue, other things? We conduct a stakeholder analysis (first used by Clark et al. for secure email systems [1]). Table 2 provides, as rows, a set of stakeholder groups and, as columns, a set of possible deliverables a quest system might provide. Deliverables are phrased to be generally favourable, so most stakeholder groups will default to being moderately in favour or, at worst, ambivalent (empty cell). Cells showcase when there is strong support, strong opposition, or internal conflict over a priority. This table helps us locate tensions between stakeholder groups.
@@
The first column describes the main function of a quest system: to enhance genuine user activity, a goal no stakeholder group opposes. Quest completions are not ‘genuine’ in the sense of economic activity—quests are completed for the sake of the quest, which is superficial. However we use ‘genuine’ to distinguish from attempts to game the quest system (which we will turn to in column 4). In the second column, again no stakeholder group opposes quests with educational value for users, while such quests are strongly preferred by exploratory users.
@@
As of writing, the blockchain space is fragmented with many competing L1s, L2s, and services for common on-chain tasks (e.g., DEXes, lending protocols, leveraged trading, etc.). Market share and user retention is a basic business goal of projects and their loyal users. Projects offering quests will typically only include quests that use the project, while quest platforms (the ‘loyalty layer’) offer a single hub for quests across multiple projects. Their niche is based on quests being spread across competitors. Occasionally projects will join quest platforms and then poach users for their own internal quest systems—called a ‘vampire attack.’
@@
The most contentious deliverable is the elimination of farming—bots, sybils, and other methods used by transactional users to automate task completion. The basic tension is that inflated completion metrics (e.g., monthly active users (MAU)) paint the picture of a fast-growing, robust user base, but its illusionary nature (‘vanity metrics’) could compromise credibility or collapse if transactional users exhaust the opportunities and move on to newer projects. Anti-farming measures are supported by DApps chosing a project to deploy on, favouring platforms with large genuine user bases. Joining them are metric platforms, which provide rankings of most popular projects, will lose credibility and relevance if it is easy for projects to use quests to game their metrics and climb the leaderboard. Loyal users receive more recognition when their contributions are not drowned out by a sea of fake engagement.
@@
4 Concluding Remarks
It is easy to see why quest systems are popular. In Web3, reaching users with a new project is challenging. Web3 does not operate with an “app store” that is easy to browse. Beyond bringing in new users, quests promise to educate users and build loyalty. However if quests are too successful, farming and bots may follow real users. Farming is a kind of “Midas touch” for projects, immediately boosting metrics and engagement. But ultimately, it forces teams into an uncomfortable dilemma: either admit that much of their previous participation came from bots—risking tension with partners, genuine users, and investors—or remain silent and appear to lose the bulk of their user base. Relying on bots to attract additional users or applications is a perilous strategy, particularly for teams that lack a clear, long-term plan to manage bot activity and know when to stop engaging with them.
Acknowledgements.
J. Clark acknowledges support for this research project from (i) the National Sciences and Engineering Research Council (NSERC), Raymond Chabot Grant Thornton, and Catallaxy Industrial Research Chair in Blockchain Technologies, (ii) the AMF (Autorité des Marchés Financiers), and (iii) NSERC through a Discovery Grant.

### v2

@@ 2.4 Thresholds
As apparent from the figure, the implementation of a threshold is associated with an immediate decline in quest completion. The purple lines follow the same trend as the grey (control) lines (dampened by not being easy tasks) until the threshold is implemented, and then plummet. One might hypothesize that bots and farmers would take a few epochs to update their scripts or acquire threshold tokens, but the association appears sticky with no future uptick. When the threshold is removed, the quests’ completion rates return to levels similar to those observed in the control group.
@@
Similarly, another quest, which had a threshold of approximately \$5 USD enforced at epoch 23, experienced a 75.8% drop in completions. After the threshold was removed at epoch 33, completions increased by 33.4%.
These examples highlight a substantial association between thresholds and lower user participation, as well as an association between the removal of thresholds and a return to higher completion rates.
2.5 Rewards with Monetary Value
@@
Figure 6: Five quests (by colour) with rewards in non-monetary points (bars without stars) and token rewards for one epoch (bars with stars). The grey line shows an average of task completion for the other quests over the same epochs. The chart shows that offering tokens (with a monetary value) as rewards is associated with higher quest completion.
@@
Consider Figure 6 which depicts 5 quests by colour, each offered over epochs 43–49 (except the yellow quest, which was first offered in epoch 46). For most epochs, the completion award was non-monetary (loyalty or experience) points, as we have analyzed in the previous section. In contrast, for one epoch (shown with a star), alongside the usual point rewards, specific quests were selected for a token reward (in the native token of the blockchain project) which had a small monetary value of  \$0.10 USD. The grey line shows the completion rate of the control quest offered across the same epochs.
@@
This was not a perfect experiment for assessing the association of monetary rewards with quest completion, as the rewards were capped at \$5000 USD, allowing a maximum of 50,000 users to claim rewards before they ran out. Further, if two quests offered monetary rewards in the same epoch, they would compete against each other until the tokens were depleted (see epoch 46 in Figure 6).
@@
Despite these defects in the data, the results are insightful: even with a minimal reward, completion rates surged. The highest increase was over 12x for a quest typically completed by around 3,500 users. With the token reward, this quest (purple in Figure 6) saw 44,196 completions. Directly after the token reward for this quest was removed, the completion rate dropped back to its original levels before the reward was introduced. Note that in the literature on loyality programs, users prefer monetary rewards as opposed to non-monetary [13].
@@
These data reveal that the introduction of a token reward is associated with a spike in completions, while the removal of the reward corresponds with a decline back to earlier levels. These spikes suggest that token rewards do not foster loyalty or habitual engagement but rather create a purely transactional relationship. Users appear to compare the cost to the reward and only engage when there is a clear opportunity for profit.
@@
In epoch 46, two quests with token rewards showed different outcomes. While the yellow quest (which had a low cost) saw a significant increase in completions, the red quest (with a cost higher than the reward) experienced no increase at all. If users were engaging with the platform out of loyalty or mastery, they would have completed the red quest as well as the yellow quest. However, this did not happen: users who claimed the reward for the yellow quest did not complete the red quest, leaving the reward for the red quest unclaimed.
3 Discussion
3.1 DQ1: Farming and Bots
We currently lack data revealing how many task completions come from genuine users as opposed to farmers (humans and bots). The Fogg behavior model [8] (B=MAP) states that a behaviour requires motivation (M), ability (A), and a prompt (P). If any element is missing, the behaviour likely will not occur. Under this model, humans and bots differ: even tiny rewards can motivate (M) bots at scale (picking up every penny), while humans aware of the tasks will likely not change (either continuing completion due to loyalty or continuing to ignore them). Bots also detect prompts (P) faster than humans, including generalized MEV bots that monitoring mempools and front-run profitable transactions [5]. This ability (A) to react instantly sets bots apart. Both farmers and bots operating from sybil addresses split tokens across many newly created addresses, each with minimal capital to cover fees, while genuine users often keep higher balances.
Putting this together, we suggest three heuristics that may help distinguish genuine users from bots: (1) If a quest alternates between unprofitable and profitable, bots will be sensitive to the exact moment it becomes marginally profitable and act; (2) if a quest requires a threshold (minimal amount of tokens), bots will be sensitive to ensuring the address meets but does not exceed the threshold amount; and (3) if a (profitable) quest is announced (especially unexpectedly), bots will be the quickest to react. We emphasize that if these heuristics were deployed in a way that harmed bots (e.g., banning rather than passive measurement), bots can adjust their behaviour to evade detection, resulting in a ‘cat and mouse’ game between farmers and quest creators. That said, we suggest new quest designs that might help identify and quantify farming activity.
@@ 3.1 DQ1: Farming and Bots
The first design observes that if costs (i.e., gas fees, service fees, buying a token or NFT as part of the quest, etc.) are slightly higher than rewards (i.e., tokens of monetary value received for quest completion), genuine users might be happy they are at least receiving a subsidy for a quest they would complete anyways (cf. [11]), while bots will avoid unprofitable tasks. When benefits are slightly higher than costs, bots will engage (including general-purpose Maximal Extractable Value (MEV) bots [5] that agnostic about the quest system). For many quest designs, profitability could vary due to external factors, such as fluctuations in gas or token values, creating the conditions for a ‘natural experiment.’ New engagement when quests become slightly profitable could be largely attributed to bots.
The second idea corresponds to the use of thresholds. If a quest can only be completed by addresses holding at least $X$ tokens for a certain period of time (days or weeks), bot farms need to divvy their tokens between their addresses. The transaction trace of this unusual activity could augment other patterns used by exiting farming-detection algorithms [7, 18].
@@
A final idea builds on the speed at which bots react. In the quest system we studied, a smart contract was randomly replenished with a fixed daily reward pool of \$500. Pre-enrolled participants with sufficient loyalty points could claim a nominal reward (\$0.01 or \$0.025, depending on tier) until the pool was depleted. Looking at Table 1, we see that depletion happened quicker each day. We also see that refill times were not always predictable. On Day 7, for instance, despite a one-hour delay in the refill time, approximately 1,500 claims were made within 2 minutes. The following day, a more drastic delay of 3 hours saw claims drop to 300, likely due to the implementation of rate-limiting and a mandatory wait period between claims. Notably, this led to a temporary spike in the time required to deplete all rewards, which rose to 2h41m. By Day 9, however, the refill occurred 4.5 hours earlier, and bots had clearly adapted, with 1,500 claims occurring within 2 minutes and the time to deplete rewards dropping to 27 minutes. Finally, following the introduction of an automated daily refill schedule on Days 10 and 11, the system reached its operational minimum, with rewards being depleted in under 30 minutes and claims stabilizing at  1,500 within the first 2 minutes.
@@
3.2 DQ2: Stakeholder Analysis
@@
A deeper question is what stakeholders actually prioritize. Is it the elimination of bots, metric boosts, revenue, other things? We conduct a stakeholder analysis. Our stakeholder analysis (Table 2) is based on our assessment as researchers of the preferences and incentives of different groups, rather than empirical data collected through surveys or interviews. This approach follows the methodology used by Clark et al. [4] for secure email systems. For space constraints, the full explanation of Table 2 is in the full version of the paper. [^2] Appendix, Full version: https://arxiv.org/abs/2501.18810
We highlight the column resilient to farming as the most contentious. DApps favor platforms with large genuine user bases, supporting anti-farming measures. Project-based stakeholders often tolerate farming or turn a blind eye. Investors accept inflated metrics when selling but seek real user data when buying, creating internal conflict. Developer teams face the same tension—wanting ground truth when selecting projects but valuing vanity metrics once committed.
4 Concluding Remarks
Many teams (based on personal conservations and social media) sense that their platform has attracted a significant number of bots. In some instances, these teams recognize that their applications or quest systems have effectively become ‘farming machines’ for bots. Farming is a kind of “Midas touch’ for projects, immediately boosting metrics and engagement. But ultimately, it forces teams into an uncomfortable dilemma: either admit that much of their previous participation came from bots—risking tension with partners, genuine users, and investors—or remain silent and appear to lose the bulk of their user base.
Acknowledgements.
The authors thank the reviewers, shepherd, and workshop chairs for pointing out many issues that greatly improved the paper and its results. J. Clark acknowledges support for this research project from NSERC, Raymond Chabot Grant Thornton, and Autorité des Marchés Financiers.
Appendix 0.A Further Stakeholder Analysis
@@
Table 2 provides, as rows, a set of stakeholder groups and, as columns, a set of possible deliverables a quest system might provide. Deliverables are phrased to be generally favourable, so most stakeholder groups will default to being moderately in favour or, at worst, ambivalent (empty cell). Cells showcase when there is strong support, strong opposition, or internal conflict over a priority. This table helps us locate tensions between stakeholder groups. The first column describes the main function of a quest system: to enhance genuine user activity, a goal no stakeholder group opposes. Quest completions are not ‘genuine’ in the sense of economic activity—quests are completed for the sake of the quest, which is superficial. However we use ‘genuine’ to distinguish from attempts to game the quest system (which we will turn to in column 4). In the second column, again no stakeholder group opposes quests with educational value for users, while such quests are strongly preferred by exploratory users.
@@
As of writing, the blockchain space is fragmented with many competing L1s, L2s, and services for common on-chain tasks (e.g., DEXes, lending protocols, leveraged trading, etc.). Market share and user retention is a basic business goal of projects and their loyal users. Projects offering quests will typically only include tasks that engage with their own platform, while quest platforms (the ‘loyalty layer’) offer a single hub for quests across multiple projects. Their niche is based on quests being spread across competitors. Occasionally projects will join quest platforms and then poach users for their own internal quest systems—called a ‘vampire attack.’
@@
The most contentious deliverable is the elimination of farming—bots, sybils, and other methods used by transactional users to automate task completion. The basic tension is that inflated completion metrics (e.g., monthly active users (MAU)) paint the picture of a fast-growing, robust user base, but its illusionary nature (‘vanity metrics’) could compromise credibility or collapse if transactional users exhaust the opportunities and move on to newer projects. Anti-farming measures are supported by DApps choosing a project to deploy on, favouring platforms with large genuine user bases. Joining them are metric platforms, which provide rankings of most popular projects, will lose credibility and relevance if it is easy for projects to use quests to game their metrics and climb the leaderboard. Loyal users receive more recognition when their contributions are not drowned out by a sea of fake engagement.